4.2
# 8- The sum of three consecutive odd integers is 81. Find the integers.

Answer by checkley71(8403)

X+(X+2)+(X+4)=81
3X+6=81
3X=81-6
3X=75
X=75/3
X=25 FIRST ODD NUMBER
X+2=25=2=27 SECOND ODD NUMBER
X+4=25=4=29 LAST ODD NUMBER
PROOF
25+27+29=81
81=81